OpenStudy (anonymous):

write the equation for the parabola whose vertex is at (3,5) and whose directix is the line y=3.

OpenStudy (anonymous):

hello is that the correct prob? is the directrix=y=3?

OpenStudy (anonymous):

that is the way its written. I have choices a. (y+5)^2=4(x-3) b. (x-3)^2=-4(y+5) c. x^2=-8(y+5) d. (x-3)^2=8(y-5) e. none of the above

OpenStudy (anonymous):

using the formula 4p(y-k)=(x-h)^2 and vertex =V(3,5) 4p(y-5)=(x-3)^2 directrix d=3=5-p p=5-3=2 therefore 4p(y-k)=(x-h)^2 8(y-5)=(x-3)^2 ans.....or 8y-40=x^2 -6x+9 x^2 -6x-8y+49=0 ans....

OpenStudy (anonymous):

the parabola opens upward